OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form.

OpenStudy (anonymous):

slope= rise/run = 9/3 = 3 y-intercept = -3 (because a point is (0,-3) so the equation is y=3x-3

OpenStudy (anonymous):

(3,6)-(0,-3) m=(-3-(6))/(0-(3)) m=3 y-6=3x-9 y=3x-3